A stone is launched vertically upwards from ground level with the initial velocity 30 m/s. Assume that the acceleration is 10 m/s².


a. Find the time taken for the stone to return to the ground again. b. Find the maximum height reached by the stone.


c. Find the velocity with which it hits the ground. d. Sketch the position-time graph and the velocity-time graph for the


motion.


P-3


e. Find the distance covered by the stone from the launch to when it


returns to earth.


1
Expert's answer
2021-06-24T16:14:43-0400

a)


"T=2\frac{30}{10}=6\ s"

b)


"H=\frac{30^2}{2(10)}=45\ m"

c)


"v_f=-30\frac{m}{s}"

d)


"s=2(45)=90"